# Continuum limit of the mobility edge and taste-degeneracy effects in high-temperature lattice QCD with staggered quarks

### 05 Dec 2023 v1

@@ 4 Conclusions
Localization of the low-lying Dirac eigenmodes in the high-temperature phase of QCD and other gauge theories [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] is closely related to the change in the confining properties of these theories [18, 32, 33, 34, 12, 26, 19, 20, 21, 22, 25, 26, 28, 29], and its study can lead to a better understanding of the mechanism behind deconfinement, and its relation with chiral symmetry restoration.
@@
In this paper we have carried out a dedicated study of the effects of taste degeneracy on the statistical properties of the staggered spectrum in high-temperature lattice QCD. We focused in particular on how these affect the numerical determination of the mobility edge, and how these effects change as the lattice spacing is reduced, or the aspect ratio is increased, with the main goal of providing a controlled continuum limit of the mobility edge.
@@
Our findings are in line with theoretical expectations, with a systematic overestimation of the mobility edge on finer lattices, where the effects of taste degeneracy become sizeable also in the bulk of the spectrum. For larger aspect ratios at fixed lattice spacing these effects are reduced and the overestimation of the mobility edge is mitigated. In the thermodynamic limit, this estimate tends to the correct value of the mobility edge, obtained independently by studying the fractal dimension of the eigenvectors. Moreover, this quantity shows a mild dependence on both volume and lattice spacing, suggesting the interchangeability of the continuum and thermodynamic limits.
@@
The most important result of this analysis is that the infinite-volume extrapolation of the mobility edge on a finer lattice is in good agreement with the mobility edge found on coarser lattices, where taste-degeneracy effects do not reach into the bulk of the spectrum and a reliable estimate can be obtained from spectral statistics already at lower aspect ratios. This shows that accurate values for the mobility edge can be obtained on relatively coarse lattices for reasonable aspect ratios using spectral statistics, which is the ideal combination from the numerical point of view.
@@
Moreover, our findings allowed us to perform the first, fully controlled extrapolation of the mobility edge to the continuum limit. This confirms the theoretical expectation [4, 35] that the mobility edge in units of the quark mass is a renormalized quantity with physical meaning, and not only a lattice artifact. It also lends support to the numerical evidence for this fact presented in Refs. [4, 29]. There it was shown that the mobility edge depends only mildly on the lattice spacing. In the light of our results, this is because the calculations of Refs. [4, 29] use relatively coarse lattices, where taste-degeneracy effects on the mobility edge are negligible. One might have wondered if the situation could have changed on finer lattices; our results show that this mild dependence would still show on finer lattices, provided one extrapolated first to the thermodynamic limit. (Incidentally, the extrapolation in $T$ shown in Ref. [4] is in good qualitative agreement with our result.)
@@
In conclusion, we have shown that one can use staggered fermions efficiently and reliably to study the mobility edge in the Dirac spectrum of high-temperature lattice gauge theories, provided that the aspect ratio is sufficiently large to avoid sizeable taste-degeneracy effects in the relevant region of the spectrum. Moreover, we have numerically demonstrated that the mobility edge in units of the quark mass is a renormalized quantity, in agreement with theoretical expectations; and that it depends only mildly on the lattice spacing.
@@ Acknowledgements.
It is a pleasure to thank Massimo D’Elia and Tamás G. Kovács for useful discussions and for a careful reading of the manuscript, and Giuseppe Clemente for help in setting up the diagonalization code. The work of CB is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033. CB also acknowledges support from the project H2020-MSCAITN-2018-813942 (EuroPLEx) and the EU Horizon 2020 research and innovation programme, STRONG-2020 project, under grant agreement No 824093. MG was partially supported by the NKFIH grant KKP-126769. Numerical calculations have been performed on the Finisterrae III cluster at CESGA (Centro de Supercomputación de Galicia). We also acknowledge the use of gauge configurations and Dirac spectra previously obtained using the MARCONI100 machine at CINECA, based on the agreement between INFN and CINECA (under projects INF19_npqcd, INF20_npqcd and INF21_npqcd).

### 12 Mar 2024 v2

@@ 4 Conclusions
Localization of the low-lying Dirac eigenmodes in the high-temperature phase of QCD and other gauge theories [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] is closely related to the change in the confining properties of these theories taking place at the transition [18, 32, 33, 34, 12, 26, 19, 20, 21, 22, 25, 26, 28, 29], and its study can lead to a better understanding of the mechanism behind deconfinement, and of its relation with chiral symmetry restoration.
@@
In this paper we have carried out a dedicated study of the effects of taste degeneracy on the statistical properties of the staggered spectrum in high-temperature lattice QCD. We focused in particular on how these affect the numerical determination of the mobility edge, and how these effects change as the lattice spacing is reduced, or the aspect ratio is increased, with the main goal of providing a controlled continuum limit of the mobility edge. To this end, we studied in detail the possibility of exchanging the order of the continuum and thermodynamic limits, arguing that it applies to the study of localization properties and the determination of the mobility edge.
@@
Our findings are in line with theoretical expectations, with a systematic overestimation of the mobility edge using spectral statistics on finer lattices, where the effects of taste degeneracy become sizeable also in the bulk of the spectrum. For larger aspect ratios at fixed lattice spacing these effects are reduced and the overestimation of the mobility edge is mitigated. In the thermodynamic limit, this estimate tends to the correct value of the mobility edge, obtained independently by studying the fractal dimension of the eigenvectors. Moreover, it agrees with finite-volume estimates obtained from spectral statistics using a suitably taste-symmetrized spectrum, for which the effects of taste degeneracy should be reduced. This supports our expectation that continuum and thermodynamic limits can be exchanged when studying the localization properties of Dirac eigenmodes.
@@
From a practical perspective, the most important result of this analysis is that the infinite-volume extrapolation of the mobility edge on a finer lattice is in good agreement with the mobility edge found on coarser lattices, where taste-degeneracy effects do not reach into the bulk of the spectrum and a reliable estimate can be obtained from spectral statistics already at lower aspect ratios. This shows that accurate values for the mobility edge can be obtained on relatively coarse lattices for reasonable aspect ratios using spectral statistics, which is the ideal combination from the numerical point of view.
@@
Moreover, our findings allowed us to perform the first, fully controlled extrapolation of the mobility edge to the continuum limit. This confirms the theoretical expectation [4, 35] that the mobility edge in units of the quark mass is a renormalized quantity with physical meaning, and not only a lattice artifact. It also lends support to the numerical evidence for this fact presented in Refs. [4, 29]. There it was shown that the mobility edge depends only mildly on the lattice spacing. In the light of our results, this is because the calculations of Refs. [4, 29] use relatively coarse lattices, where taste-degeneracy effects on estimates of the mobility edge based on spectral statistics are negligible. One might have wondered if the situation could have changed on finer lattices; our results show that this mild dependence would still show on finer lattices, provided one extrapolated first to the thermodynamic limit. (Incidentally, the extrapolation in $T$ shown in Ref. [4] is in good qualitative agreement with our result.)
@@
In conclusion, we have shown that one can use staggered fermions efficiently and reliably to study the mobility edge in the Dirac spectrum of high-temperature lattice gauge theories, provided that the aspect ratio is sufficiently large to avoid sizeable taste-degeneracy effects in the relevant region of the spectrum. The possibility to bypass these effects through this procedure is due to the possibility of exchanging the continuum and thermodynamic limits when studying localization properties. Moreover, we have numerically demonstrated that the mobility edge in units of the quark mass is a renormalized quantity, in agreement with theoretical expectations; and that it depends only mildly on the lattice spacing.
@@ Acknowledgements.
It is a pleasure to thank Massimo D’Elia and Tamás G. Kovács for useful discussions and for a careful reading of the manuscript, and Giuseppe Clemente for help in setting up the diagonalization code. The work of C. B. is supported by the Spanish Research Agency (Agencia Estatal de Investigación) through the grant IFT Centro de Excelencia Severo Ochoa CEX2020-001007-S and, partially, by grant PID2021-127526NB-I00, both funded by MCIN/AEI/10.13039/501100011033. C. B. also acknowledges support from the project H2020-MSCAITN-2018-813942 (EuroPLEx) and the EU Horizon 2020 research and innovation programme, STRONG-2020 project, under grant agreement No 824093. M. G. was partially supported by the NKFIH grant KKP-126769, and by the NKFIH excellence grant TKP2021-NKTA-64. Numerical calculations have been performed on the Finisterrae III cluster at CESGA (Centro de Supercomputación de Galicia). We also acknowledge the use of gauge configurations and Dirac spectra previously obtained using the Marconi100 machine at Cineca, based on the agreement between INFN and Cineca (under projects INF19_npqcd, INF20_npqcd and INF21_npqcd).